OpenStudy (anonymous):

A missile is launched from the ground with a speed of 100 m/s at an initial angle of 45 degrees to the ground. In what direction is it moving just 2 seconds before hitting the ground?

OpenStudy (anonymous):

calculate the total time first: In the y-direction: Dy = 0 m, V1y = 70.71 m/s, a = 9.8 m/s^2. Take one of the kinematic equations to find time, \[dy = v1yt + 1/2at^2\] \[0 = 70.71t - 4.9t^2\] use quadratic formula to solve for t, t = 14.43 s 14.43 s - 2 s= 12.43 s (because the question says 2 seconds before hitting the ground) Therefore, we have Dy = 0 m, V1y = 70.71 m/s, a = -9.8 m/s^2, t = 12.43 Now, find the vertical component of the final velocity. Vf = V2y + V2x Take one of the kinematic equations again to find V2y. \[v2y = v1y + at\] \[v2y = 70.71 - 9.8(12.43)\] V2y = -51.104 m/s (the negative sign indicates the velocity in the down direction)  Find the V2x component. V2x = 100cos45 = 70.71 m/s Use the tangent ratio to find the angle. \[\tan \theta = opp/adj\] \[\tan \theta = 51.104/70.71\] \[\tan^{-1} (51.104/70.71) = 35.8 \] Therefore, the missile is moving at a direction of 35.8 degrees below the horizontal 2 seconds before hitting the ground.